OpenStudy (anonymous):

Which of the following statements follows from (x - 3)^2 = 5? x^2 + 9 = 5 x - 3 = ± x - 3 = ±25

OpenStudy (anonymous):

x-3=\[\pm \sqrt{5}\]

OpenStudy (anonymous):

thats the second answer

OpenStudy (anonymous):

the missing one!

OpenStudy (ajprincess):

vat's ur answer?

OpenStudy (binary3i):

(x-3)^2 = 5\[x -3 = \pm \sqrt{5}\]

OpenStudy (anonymous):

no i needed to add the end of it it was missing prt of it

OpenStudy (anonymous):

ok thanks